find the area of a rectangle with length 16 inscribed in a circle with radius 10
Answer by rwm(914)

the diagonal(hypotenuse) is 20
2*r=20
a^2+b^2=c^2
a=16
c=20
a*b=d
find b and multiply it by 16 to get the area
b=12
d=16*12=192